OpenStudy (anonymous):

What is the additive inverse of 4x?

OpenStudy (turingtest):

the additive inverse of n is m such that n + m = 0 what can you add to the above to get zero?

OpenStudy (anonymous):

so would it b x - 4

OpenStudy (turingtest):

no, because 4x+(x-4)=4x+x-4=5x-4 which is not zero...

OpenStudy (anonymous):

ohhh -4x

OpenStudy (turingtest):

yep :) nice job!